Standing Committee Meeting Report PASSCAL

March 2014

Action items from PASSCAL Standing Committee
Spring 2014 meeting - Washington, DC - IRIS HQ

1.  Gridley, K Anderson:  Post agenda and action items to the web for the most recent PASC meetings

Status:  K Anderson will take this on with Justin - Action Item open

2. K Anderson:  Receive report from the January 2014 Active Source meeting.  Follow up with PASC call to review the report.  Then to be distributed through CoCom/BoD and community.

Status:  Report draft not yet complete.  K Anderson will distribute current draft to PASC.  This will be discussed at Fall PASC meeting - Action item open

3. K Anderson:  Include PASC members on distribution of facility quarterly reports

Status:  Accomplished - Reports for last 2 quarters distributed to PASC.  All reports will be posted on PASC webpage and future reports will be posted and sent to PASC members after PM approval - Action item closed

4. K Anderson:  Query Andy F about how to update PASC webpage.

Status:  K Anderson will take this on and build a website similar to GSN SC page with access to old meeting information, current meeting documents, and historical documents (e.g. distribution point for PIC quarterly reports) -  Action item closed

5. K Anderson:  Gridley will get list of contacts from SEG to Ted Clarke.

Status:  Gridley no longer working for IRIS, so this is no longer required - Action item closed

6. Woodward/Busby:  Provide plan and equipment numbers for CEUSN and equipment availability

Status:  Woodward will distribute current plan and discuss at Fall PASC meeting - Action item closed

7. K Anderson/Moran/SC:  Come up with a plan to absorb TA systems into the PASSCAL equipment pool, along with any impacts.

Status:  Anderson and Beaudoin will draft plan - Action item remains open

8. Beaudoin:  Provide data from TA for turn around / uses vs. repair.

Status  Completed and part of Sustainability part I report - Action item closed

9. Beaudoin: Send UTEP Texans from the PIC to UTEP

Status:  Completed - Action item closed

10. Beaudoin:  Provide clarity on data policy to PIC staff

Status:  Completed - Action item closed

11. Moran:  Update and discuss outstanding FA merger issues with BoD (data services, usage policy for experiments outside US)

Status:  Seth will review his notes on these topics

12. Moran:  SRL remembrance article on Jim Fowler

Status:  Completed - Action item closed

13. Woodward:  Status of Sweetwater dataset

Status:  To be presented at PASC in IS section - Action item closed

14. K Anderson:  Post all PASC PowerPoint presentations on PASC webpage

Status:  This has been accomplished for Oct 14 meeting - Action item closed

15. Beaudoin:  Present status of direct burial results at IRIS workshop

Status:  KatyLiz made presentation in a poster at workshop - Action item closed

16. K Anderson:  Get PNSC action items to PASC

Status:  K Anderson will work with Nyblade on this.  This will become part of PNSC SOP.  K Anderson will make Polar related status reports available to PASC, including SAGE budget requests, GLISN, GEOICE, and supplemental proposals - Action item remains open

17. K Anderson / PNSC: Vet recommendations for PNSC rotation (Nettles coming off)

Status:  Completed - Paul Winberry has now replaced Nettles on PNSC - Action item closed

18. Beaudoin:  Report on number of non-NSF funded experiments supported by PASSCAL

Status:  Completed and is included in the Sustainability report - Action item closed

19. Beaudoin:  Calculate average number of instruments per experiment

Status:  Completed and is included in the Sustainability report - Action item closed

20.  Moran:  Set up conference call to discuss recommendations from PASWG

Status: Completed - Action item closed

21.  Okaya:  Develop science highlights and bullets for presentation to BoD and CoCom to showcase PASSCAL science accomplishments

Status:  Completed and presented to BoD and CoCom - Action item closed

22.  Beaudoin / Okaya:  Contact Katie K. about TSP vs PH5 (Gridley)

Status:  Laptop was configured and sent to Keranen for testing both software packages.  The laptop was returned and it is not clear if the software packages were exercised.  Bruce will follow up with Keranen.  This discussion will continue as part of the tasking of the ASWG (Moran/Wagner/Hole/Miller/Keranen/K Anderson) - Action item is ongoing

23.  Beaudoin:  Collate recent RAMP examples including impact

Status:  Information will be presented as part of PIC Ops presentation.  In brief, the RAMP pool is 10 6-channel DAS, T120 (10) and episensors (10). However, there are 6 currently active RAMP's in the field now [Chicago (1 BB), Greeley (6 SP), Chihuahua (7 SP), Arizona (6 SP), and 2 in Napa (7 BB + 7 Accel + 270 Texans)]. So essentially we have WAY more than the standard RAMP out and the instruments are quite a bit more diverse - Action item closed

Action Items from discussion with Greg Anderson

24.  Moran:  Summarize conversation between Greg and PASC from Spring meeting and send to Board Chair (Anne M) and PASSCAL Board Liaison (Geoff A)

Status:  Completed - Action item closed

25.  Moran:  Send Greg revised recommendations from PASC on the merger of FA instruments into PASSCAL

Status:  Completed - Action item closed

26.  Moran:  Send one-pagers to Greg clarifying current data policies and current "satisfaction" level with current policies

Status:  Unsure if this was completed, Moran will verify - Action item remains open

27.  Okaya/K Anderson/Beaudon/Sumy:  Follow up with SRL editor on experiment description articles in SRL

Status:  Zhigang Peng (SRL Editor) has been contacted and is interested in working with IRIS on a template for experiment descriptions/introductions that can be used as citable references to those using the experiment data.  Sweetwater will be the initial example and the template will then be distributed.  Okaya also mentions a Nature publication for Scientific Dataset ([url=][/url]) - Action item closed

28.  Moran/K Anderson:  Establish PASC subcommittee and address specific issue of sensor decommission

Status:  Intent was to draft charges to Instrumentation subcommittee on establishing thresholds for performance and decommissioning standards.  Anderson and Beaudoin will try to draft something before PASC meeting - Action item remains open

29.  Moran/PASWG:  Develop scope management process and send to BoD

Status:  Completed - Accomplished by Moran and PASC.  Document drafted and distributed to BoD for May meeting - Action item closed

30.  Hole:  Rewrite New Technology report summary and submit to PASC for CoCom

Status:  Summary of executive summary summarized and sent to CoCom and BoD - Action item closed

31.  Moran:  Plan PASSCAL SIG at IRIS workshop

Status: Completed - SIG held, focus was on sustainability - Action item closed

32.  K Anderson:  Organize a call to wrap up Strategic Planning document

Status:  This did not happen.  We will make plans to complete the strategic plan at the Fall 2014 PASC meeting - Action item remains open

Public Action Items from Executive Session:

33. K Anderson:  Work with Andy Frassetto on preparing a "PASSCAL 101" webinar for ECI community

Status:  Follow up with Andy - Action item open

34. Okaya/K Anderson/Wagner:  Represent PASSCAL community as liaison at ECI breakfast at AGU

Status:  Completed - Action item closed

35. Moran:  Discuss use of TA instruments from lower 48 to offset aging instruments in PASSCAL pool

Status:  Many conversations held.  This will be presented at IS section of PASC meeting.  Also a topic for discussion at ISSC meeting - Action item closed